OpenStudy (anonymous):

what is (-m) to -3 exponet times n, if m=2 and n=-24? A.3 B.-3 C.4 D.-4

hartnn (hartnn):

so, \(\Large (-2)^{-3} \times (-24)\)

hartnn (hartnn):

\(\Large (-2)^{-3} = \dfrac{1}{(-2)^3}= ... \) ??

hartnn (hartnn):

can you find (-2)^3 ?

hartnn (hartnn):

(-2)^3 = (-2)(-2)(-2) = ...

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

YES so, \((-2)^{-3} \times 24 = \dfrac{-24}{-8} =... ?\)

OpenStudy (anonymous):

3

OpenStudy (anonymous):

thank you so much

hartnn (hartnn):

yes, 3 :) welcome ^_^